OpenStudy (anonymous):

Divide and simplify b+5/b+3 divided by ab+45/b-7

OpenStudy (anonymous):

Hi Ms Juanita...ur on a roll.....Divide and simplify b+5/b+3 divided by ab+45/b-7

OpenStudy (lgbasallote):

is that ab + 45? or 9b + 45? the latter makes more sense

OpenStudy (anonymous):

Im sorry its 9b+45

OpenStudy (anonymous):

agreed (b-7) /9(b+3)

OpenStudy (lgbasallote):

get the reciprocal... b+5/b+3 x b-7/9b+45 b+5/b+3 x b-7/9(b+5) b-7/9(b+3)

OpenStudy (anonymous):

y+4/y=-5

OpenStudy (anonymous):

im assuming tha this is solving for y only ? y = -2/3

OpenStudy (lgbasallote):

is that another question?

OpenStudy (anonymous):

yes it is juanita

OpenStudy (anonymous):

it also says use a comma to separate answers

OpenStudy (anonymous):

confirm that you wrote the prolem in its entirety

OpenStudy (anonymous):

yes it is y+4/y=-5

OpenStudy (anonymous):

The options it gives are solutions or no solution

OpenStudy (anonymous):

v+2/v+v/v+2

OpenStudy (anonymous):

oh - 2nd is solutions ar -2, -1 for the first solutions are -4,-1

OpenStudy (anonymous):

v+2/v+v/v+2

OpenStudy (anonymous):

the answer is solutions are -2, -1

OpenStudy (anonymous):

for which one this one; v+2/v+v/v+2

OpenStudy (anonymous):

the answer is solutions are -2, -1

OpenStudy (anonymous):

-2, -1 for v+2/v+v/v+2

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okaqy thanks ...u been a big help....im in the library and they are closing, i will resume this tomorrow...I hope that you are on......